OpenStudy (anonymous):

find the derivative of f(x)=1/x(x+1)

OpenStudy (lgbasallote):

\[f(x) = \frac{1}{x(x+1)}\] is that the question?

OpenStudy (lgbasallote):

are you familiar with partial fractions?

OpenStudy (anonymous):

that is the question

OpenStudy (anonymous):

@lgbasallote just do (1/v)'

OpenStudy (lgbasallote):

if mariaad knows partial fractions it would be easier...

OpenStudy (anonymous):

i got -2x-1/(x^2)(x+1)^2

OpenStudy (anonymous):

x^2(x+1)^2 as the denom

OpenStudy (lgbasallote):

im thinking the denominator is wrong..

OpenStudy (lgbasallote):

you should put it as \[\frac{1}{x^2 + x}\] then you can do the derivatives

OpenStudy (anonymous):

so what is the denominator

OpenStudy (lgbasallote):

it would be \[\frac{ -(2x +1)}{(x^2+x)^2} \implies - \frac{2x - 1}{(x^2+x)^2}\]

OpenStudy (anonymous):

sorry i need it for an online homework that is due at midnight

OpenStudy (lgbasallote):

so the denominator is your only mistake

OpenStudy (anonymous):

-2x - 1/((x2) + x)2.

OpenStudy (anonymous):

when i plug that in it says it is wrong

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

hmm?

OpenStudy (lgbasallote):

try expanding the denominator?

OpenStudy (anonymous):

nothing is working ! ahh

OpenStudy (lgbasallote):

hmm wait....it seems you were right the first time haha sorry

OpenStudy (lgbasallote):

but i dont think i was wrong....guess they just prefer it written the way you did..

OpenStudy (anonymous):

well either way when i type in that answer it says it is wrong i was wondering if maybe theres another way to write it maybe the parentheses are wrong or something

OpenStudy (lgbasallote):

you can write it the way you did before \[\frac{-2x-1}{x^2(x+1)^2}\]

OpenStudy (lgbasallote):

thye're just equal

OpenStudy (anonymous):

i know that I am saying that when I type it into the program it cant have the exponent like that it has to be x^2 for example so when i attempted to put that in the form it says it is wrong ...i cant figure out how else to write it